OpenStudy (anonymous):

Is (1, -3) a solution to 3y = 2x + 1?

OpenStudy (anonymous):

it is if after putting 1 = x and -3 = y, you get the same number on both sides which in this case, it looks like it will not be

OpenStudy (ash2326):

Let's check if (1, -3) is a solution \[3y=2x+1\] put x=1 y=-3 \[3\times -3=2\times 1+1\] Check if both sides are equal, if they are then it's a solution otherwise not